OpenStudy (prettygirl_shynice):

which equation can be used determine what number ,n,is 40% of 64 a.(40) (64)=n b.40÷ 64=n c.0.40 (64)=n d.64n=0.40 will medal +fan

OpenStudy (mathmale):

Why not do this in the same way? Find your own answer first, and then compare your procedure to the four answer choices.

OpenStudy (prettygirl_shynice):

ok so i still cross multiply

OpenStudy (mathmale):

n is 40% of 64. 1. Convert 40% to 40/100=0.40 2. Then write the equation (different from before) n = 0.40 * 64.

OpenStudy (mathmale):

No cross mult. here (altho perhaps someone else might use cross mult).

OpenStudy (mathmale):

What is the value of n?

OpenStudy (prettygirl_shynice):

idk

OpenStudy (prettygirl_shynice):

2.25.6

OpenStudy (prettygirl_shynice):

i think c @mathmale

OpenStudy (prettygirl_shynice):

@jigglypuff314

OpenStudy (prettygirl_shynice):

@xapproachesinfinity am ii doing it right

OpenStudy (xapproachesinfinity):

you got the answer already

OpenStudy (prettygirl_shynice):

@mathmale did do great

OpenStudy (mathmale):

Thank you! about 20 min. ago, I wrote: "n is 40% of 64." This translates to "n=0.40(64). Which one of the answer choices, if any, is closest to that?

OpenStudy (prettygirl_shynice):

i thought c

OpenStudy (prettygirl_shynice):

and yea